Please help.

usinessCourse E Menu İ QUESTION! Not complete Marked out of 9.00 F Flag question Determine and Use Overhead Rate The following selectedledger accounts of Cameron Company are for February (the second month of ts accounting year) Fir Materials Inventory Feb. 1 balance 31,500 February credits 113,000 February debits 104,000 Manufacturing Overhead February debits 137,200 Feb. 1 balance 11,600 February credits 136,350 Work in Process Inventory Feb. 1 balance February debits: Direct material Direct Labor Man. overhead 22,400 February credits 345,000 95,000 151,500 136,350 Wages Payable February debits 193,500 Feb. 1 balance 45,000 February credits177,000 Finished Goods Inventory Feb. 1 balance 76,500 February credits 383,700 February debits345,000 a Determine the am O Type here to search 2

Blackboar Homework https//mybusinesscourse.com/platfom/mod/quiz /attempt.phpřattempt 2300340&cmid-164029 usinessCourse E Menu a. Determine the amount of indirect material requisitioned for production during February b. How much indirect labor cost was apparently incurred during February? c. Calculate the manufacturing overhead rate based on direct labor cost. d. Was manufact Manufacturing overhead was e. Was manufacturing overhead for the first two months of the year under-or uring overhead for February under-or overapplied, and by what amount? e by $ for February overapplied, and by what amount? Manufacturing overhead was f. What is the cost of production completed in February? for the first two months. g. What is the cost of goods sold in February? Check
Answer #1
a
Amount of Indirect materials = Materials credit-Direct materials used = 113000-95000= 18000
b
Indirect labor cost = 177000-151500= $25500
c
Manufacturing overhead rate = 136350/151500= 90%
d
Manufacturing overhead under-applied for February = 137200-136350= $850
e
Manufacturing overhead over-applied for first two months =11600+136350-137200= $10750
f
Cost of Production completed in February = $345000
g
Cost of goods sold in February = $383700